Prahm Corp. wants to raise $4.1 million via a rights offering. The company currently has 470,000 shares of common stock outstanding that sell for $42 per share. Its underwriter has set a subscription price of $17 per share and will charge the company a spread of 5 percent.

If you currently own 5,000 shares of stock in the company and decide not to participate in the rights offering, how much money can you get by selling your rights?

I tried many answers like( 5000*8.767, 5000*8.77, 5000*8.8, 5000*9) all are wrong.

5 Prahm Corp. wants to raise $4.1 million via a rights offering. The company currently has 470,000 shares of common stock outstanding that sell for $42 per share. Its underwriter has set a subscription price of $17 per share and will charge the company a spread of 5 percent. 20 polnts If you currently own 5,000 shares of stock in the company and decide not to participate in the rights offering, how much money can you get by selling your rights? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) eBook Print References Your proceeds from sale of rights 43,850.00
Answer #1
Net proceeds to firm = $17 (1-.05) = $16.15
New share price = =4100000/16.15
253869.969
No of right shares per share =470000/253870
1.851341238
Price Exright =($17 + 1.85134(42)) / (1+1.8513)
33.23219639
Right value =42-33.23
8.767803611
Sale Proceeds =5000*8.7678
43839.01806